# Reducing Pilots in Channel Estimation With Predictive Foundation Models

### v1

@@ Abstract
Accurate channel state information (CSI) acquisition is essential for modern wireless systems, which becomes increasingly difficult under large antenna arrays, strict pilot overhead constraints, and diverse deployment environments. Existing artificial intelligence-based solutions often lack robustness and fail to generalize across scenarios. To address this limitation, this paper introduces a predictive-foundation-model-based channel estimation framework that enables accurate, low-overhead, and generalizable CSI acquisition. The proposed framework employs a predictive foundation model trained on large-scale cross-domain CSI data to extract universal channel representations and provide predictive priors with strong cross-scenario transferability. A pilot processing network based on a vision transformer architecture is further designed to capture spatial, temporal, and frequency correlations from pilot observations. An efficient fusion mechanism integrates predictive priors with real-time measurements, enabling reliable CSI reconstruction even under sparse or noisy conditions. Extensive evaluations across diverse configurations demonstrate that the proposed estimator significantly outperforms both classical and data-driven baselines in accuracy, robustness, and generalization capability.
@@ I Introduction
Accurate channel state information (CSI) plays a pivotal role in the design of wireless communication systems. The effectiveness of core technologies, including massive multiple-input multiple-output (MIMO) and orthogonal frequency division multiplexing (OFDM), critically depends on accurate CSI acquisition [1]. To meet the growing demand for ubiquitous connectivity and intelligent communications, wireless systems have continually scaled their antenna array dimensions, a trend expected to accelerate in sixth-generation (6G) networks [2]. In current 5G systems, the number of antenna ports typically ranges from 16 to 128, whereas 6G is projected to scale this number to 512 and beyond [3, 4]. This rapid increase in system dimensionality imposes a substantial burden on channel acquisition because pilot signals must be transmitted over increasingly dense time-frequency grids. Therefore, the development of channel estimation schemes that reduce pilot overhead while preserving high estimation accuracy has become a central research problem.
@@ III-A Overall Framework and Workflow
Applying the proposed framework to the channel estimation problem requires resolving several practical issues. Specifically, the formulation in (5) presumes that the PFM has access to perfect historical CSI, which is unattainable in practice and thus requires an alternative construction of the historical context. To this end, we develop a slot-based workflow as depicted in Fig. 1(b). The key design principle is a recursive estimation process where the output of slot (i1)(i-1) serves as the historical input for slot $i$. To formalize this recursive workflow and facilitate NN processing, let $\mathbf{Z}^{(i)}\in\mathbb{R}^{2N_{\rm t}K\times T}$ and $\hat{\mathbf{Z}}^{(i)}\in\mathbb{R}^{2N_{\rm t}K\times T}$ denote the real-valued tensor representations of $\mathbf{H}^{(i)}$ and $\hat{\mathbf{H}}^{(i)}$, respectively. Furthermore, let $\mathbf{Z}_{\rm his}^{(i)}$ denote the historical input fed into the PFM at slot $i$. Considering the transmission over multiple slots, as shown in Fig. 1(b), we have the following two cases.
@@ III-B2 Input Projection and Positional Encoding
where $\operatorname{PE}(\cdot)$ denotes the deterministic sinusoidal positional encoding function, and $\mathbf{m}\in\mathbb{R}^{N_{\rm pat}\times 1}$ is the patch index vector. The resulting tensor $\mathbf{E}\in\mathbb{R}^{N_{\rm pat}\times L_{\rm pat}}$ serves as the input to the transformer.
@@ III-C1 Input Feature Preparation
We first derive an initial coarse estimate, $\hat{\mathbf{H}}_{\rm P}\in\mathbb{R}^{2N_{\rm t}\times K\times T}$, based on the least squares (LS) method and interpolation, where the real and imaginary parts are concatenated along the antenna dimension. This estimate serves as a noisy approximation of the true CSI. To further suppress the noise and leverage its zero-mean property, a despreading operation [37] is adopted to average the estimate across $R_{\rm T}$ and $R_{\rm F}$ neighboring REs in the time and frequency domains, respectively, thereby enhancing feature consistency. The feature maps before and after the despreading operation are concatenated along the time dimension, resulting in a feature tensor of size $2N_{\rm t}\times K\times L_{\rm f}$, where $L_{\rm f}=T+T/R_{\rm T}$. To capture spatial correlations, the features within this tensor are grouped by subcarrier. Specifically, the spatial-time domain feature on each subcarrier is partitioned into $2N_{\rm t}$ patches of length $L_{\rm f}$, which are then linearly projected onto a latent space of dimension $d_{\rm m}$ via a fully-connected (FC) layer. This yields the input embeddings for the subsequent transformer encoder.
@@ IV-A1 Simulation Settings
The system configurations and simulation settings for channel dataset generation are summarized in Table I. The simulation parameters are randomly sampled from the specified ranges in the table to ensure diverse channel conditions. Each slot consists of $T=14$ OFDM symbols. A total of 8 RBs are allocated, resulting in $K=96$ subcarriers per slot. To emulate realistic propagation environments, the tapped delay line (TDL) channel model compliant with the 3GPP specifications [40] is employed for dataset generation. Sparse pilot configurations employing FDM are adopted in the simulations, as depicted in the top-left corner of Fig. 1(a). The “2P” and “4P” patterns denote two different allocations of pilot OFDM symbols under $N_{\rm t}=4$, with $T_{\rm p}=2$ and $T_{\rm p}=4$, respectively, used for different mobility scenarios. These pilot arrangements follow the demodulation RS (DMRS) configuration type 2 defined in 5G NR [41]. In total, 768,000 channel trajectories are generated, with each trajectory comprising 10 consecutive slots. This results in a dataset of 7,680,000 samples, each comprising a triplet of previous and current channel realizations. The dataset is partitioned into training, validation, and test sets using an 8:1:1 split ratio.
@@ IV-A3 Baselines
(a) 30 km/h
(b) 90 km/h
(c) 300 km/h
@@ IV-B Prediction Performance of PFM
To illustrate the superiority of the PFM and the motivation for its adoption, we compare this model with the state-of-the-art channel prediction method, LLM4CP [26], as shown in Fig. 5. For a fair comparison, the channel datasets released in [26] are used for training, fine-tuning, and performance evaluation of the PFM. To explore effective strategies for leveraging the PFM, four training paradigms are examined:
@@
The comparative results, depicted in Fig. 5, lead to several important observations. First, the finetuned PFM substantially outperforms the pre-trained version, indicating that adaptation using wireless channel data effectively enhances model performance in channel acquisition. Second, when comparing models trained from scratch with those finetuned from pre-trained initialization, the results highlight the advantages of leveraging pre-trained weights from the general time-series foundation model [33]. Finally, although freezing the backbone during finetuning achieves competitive performance, updating the backbone parameters yields additional gains, underscoring the importance of jointly refining the entire model to adapt the PFM to wireless channel-specific tasks. These observations substantiate the training strategy proposed in Section III-E for effectively adapting the PFM to wireless channel estimation.
@@ IV-C1 In-Distribution Performance
We first evaluate the in-distribution performance of the proposed PFM-aided channel estimation scheme (PFM-aided CE) compared to baselines and explore the model’s ability to handle varying configurations, including user speeds and pilot configurations. The channel estimation NMSE performance across different user velocities is presented in Fig. 6. As shown in the figure, PFM-aided CE outperforms all baselines by 1–5 dB in NMSE across all velocities and SNRs under both the “2P” and “4P” pilot configurations. This gain primarily stems from the model’s ability to effectively exploit temporal knowledge from previously estimated CSI. Specifically, at moderate speeds of 30 km/h and 90 km/h, the PFM-aided CE with the sparse “2P” configuration even achieves performance comparable to LMMSE using the denser “4P” configuration, especially in the low-SNR regime. At high mobility (300 km/h), all compared approaches experience performance degradation as expected, while the proposed estimator still delivers the best performance and achieves even more pronounced gains. These results confirm the ability of the model to handle channel variations with different time granularities.
@@
Next, the NMSE performance with respect to the slot index is demonstrated in Fig. 7, where a moderate speed of 90 km/h and SNR of 15 dB are considered. This evaluation aims to verify the effectiveness of the proposed workflow. The results showcase that the proposed estimator maintains stable performance across multiple slots for slot indices $i\geq 2$, clearly outperforming the first slot, which relies solely on the pilot observations for estimation. This observation highlights the effectiveness of leveraging historical channel information and the PFM’s predictive abilities. Particularly, PFM-aided CE with “2P” shows virtually the same estimation accuracy as LMMSE with “4P”, despite using only 50% of the pilot overhead. These results underscore the potential of the proposed scheme in reducing pilot overhead and enabling the design of more efficient RS configurations.
@@ IV-C2 Zero-Shot Generalization Study
We further investigate the NMSE performance under an unseen environment using the line-of-sight TDL-D profile, as shown in Fig. 10. This setting introduces a substantial distributional shift in channel correlation and delay structure, as the model was trained exclusively on non-line-of-sight scenarios (TDL-A/B/C). Despite this drastic mismatch, the proposed PFM-aided CE maintains competitive performance, remarkably surpassing the LMMSE, CNN, and ViT-based estimators. These results underscore the model’s zero-shot generalization ability, enabling adaptation to structurally different channel environments without finetuning. This capability stems from the time-series forecasting knowledge inherited from the pre-trained PFM. This large-capacity model has been pre-trained across a large-scale time-series corpus, thereby generalizing effectively to out-of-distribution wireless channel sequences, an ability inaccessible to conventional statistical or small-scale AI models.
@@ IV-D End-to-End Bit Error Rate (BER) Performance
Figure 11: End-to-end BER performance of the communication link employing the proposed channel estimator.
To assess the practical viability of the proposed channel estimator in a communication link, we evaluate its impact on end-to-end BER. We consider a multiple-input single-output OFDM system, and the channel settings follow those presented in Table I. The 5G NR low-density parity-check (LDPC) code with a rate of 466/1024 [43] and a codeword size of 7296 bits is employed to encode the message bits for all data REs. The encoded bits are then modulated using 64-ary quadrature amplitude modulation (64-QAM) and mapped onto the data REs over the resource grid. The data stream is precoded for transmission across $N_{\rm t}=4$ antennas. The transmitter is supposed to have perfect CSI, and the digital precoder on the $(k,t)$-th RE, $\mathbf{v}_{k,t}\in\mathbb{C}^{4\times 1}$, is obtained based on match filtering [42]:
$\mathbf{v}_{k,t}=\frac{{\mathbf{h}}_{k,t}}{\|{\mathbf{h}}_{k,t}\|_{2}},$ (18)
where ${\mathbf{h}}_{k,t}$ represents the CSI on the $(k,t)$-th RE. Hence, the received signal at the $(k,t)$-th RE is given by
@@
where $x_{k,t}$ and $n_{k,t}$ denote the transmitted symbol and the AWGN with variance $\sigma^{2}$ at the $(k,t)$-th RE, respectively.
At the receiver side, the received signals are processed using an LMMSE equalizer based on the estimation of the equivalent channel $\mathbf{h}_{k,t}^{\rm H}\mathbf{v}_{k,t}$. The equalized signals are then soft-demapped and delivered to the LDPC decoder to recover the transmitted message bits. The BER is evaluated after $10^{4}$ codewords are transmitted.
The BER results obtained using different channel estimators under various mobility conditions are presented in Fig. 11. For reference, the BER performance with perfect CSI at the receiver is included as an upper bound. Fig. 11(a) presents the results for a user speed of 90 km/h with the “2P” pilot pattern. As shown in the figure, the proposed estimator enables a substantial BER reduction in the high-SNR regime, surpassing the LMMSE-, CNN-, and ViT-based estimators by over 3 dB and closely approaching the upper bound. These results highlight the critical role of the accurate channel acquisition achieved by the proposed PFM-aided CE in enhancing the overall system performance with only limited pilot overhead. Fig. 11(b) further evaluates the estimators at a higher user speed of 300 km/h using the “4P” pilot pattern. Although the performance gap between the proposed scheme and the upper bound becomes larger due to increased channel dynamics, the proposed method continues to outperform all baselines by a significant margin, demonstrating strong robustness under high-mobility conditions.

### v2

@@ Abstract
Accurate channel state information (CSI) acquisition is essential for modern wireless systems, which becomes increasingly difficult under large antenna arrays, strict pilot overhead constraints, and diverse deployment environments. Existing artificial intelligence-based solutions often lack robustness and fail to generalize across scenarios. To address this limitation, this paper introduces a predictive-foundation-model-based channel estimation framework that enables accurate, low-overhead, and generalizable CSI acquisition. The proposed framework employs a predictive foundation model trained on large-scale cross-domain data to extract universal channel representations and provide predictive priors with strong cross-scenario transferability. A pilot processing network based on a vision transformer architecture is further designed to capture spatial, temporal, and frequency correlations from pilot observations. An efficient fusion mechanism integrates predictive priors with real-time measurements, enabling reliable CSI reconstruction even under sparse or noisy conditions. Extensive evaluations across diverse configurations demonstrate that the proposed estimator significantly outperforms both classical and data-driven baselines in accuracy, robustness, and generalization capability.
@@ I Introduction
Accurate channel state information (CSI) plays a pivotal role in the design of wireless communication systems. The effectiveness of core technologies, including massive multiple-input multiple-output (MIMO) and orthogonal frequency division multiplexing (OFDM), critically depends on accurate CSI acquisition [1]. To meet the growing demand for ubiquitous connectivity and intelligent communications, wireless systems have continually scaled their antenna array dimensions, a trend expected to accelerate in sixth-generation (6G) networks [2]. In current fifth-generation (5G) systems, the number of antenna ports typically ranges from 16 to 128, whereas 6G is projected to scale this number to 512 and beyond [3, 4]. This rapid increase in system dimensionality imposes a substantial burden on channel acquisition because pilot signals must be transmitted over increasingly dense time-frequency grids. Therefore, the development of channel estimation schemes that reduce pilot overhead while preserving high estimation accuracy has become a central research problem.
@@ III-A Overall Framework and Workflow
Applying the proposed framework to the channel estimation problem requires resolving several practical issues. Specifically, the formulation in (5) presumes that the PFM has access to perfect historical CSI, which is unattainable in practice and thus requires an alternative construction of the historical context. To this end, we develop a slot-based workflow as depicted in Fig. 1(b). The key design principle is a recursive estimation process where the output of slot (i1)(i-1) serves as the historical input for slot $i$. To formalize this recursive workflow and facilitate NN processing, let $\mathbf{Z}^{(i)}\in\mathbb{R}^{2N_{\rm t}K\times T}$ and $\hat{\mathbf{Z}}^{(i)}\in\mathbb{R}^{2N_{\rm t}K\times T}$ denote the real-valued tensor representations of $\mathbf{H}^{(i)}$ and $\hat{\mathbf{H}}^{(i)}$, respectively, which are obtained by concatenating their real and imaginary components along the antenna dimension. Furthermore, let $\mathbf{Z}_{\rm his}^{(i)}$ denote the historical input fed into the PFM at slot $i$. Considering the transmission over multiple slots, as shown in Fig. 1(b), we have the following two cases.
@@ III-B2 Input Projection and Positional Encoding
where $\operatorname{PE}(\cdot)$ denotes the deterministic sinusoidal positional encoding function, and $\mathbf{m}\in\mathbb{R}^{N_{\rm pat}\times 1}$ is the patch index vector. The resulting tensor $\mathbf{E}\in\mathbb{R}^{N_{\rm pat}\times d}$ serves as the input to the transformer.
@@ III-B4 Output Projection and Denormalization
Remark 2
The cross-domain transferability of the PFM from macroscopic data, e.g., weather and traffic, to microscopic channel fading is grounded in the mathematical isomorphism of temporal dynamics. Despite vastly different physical time scales, both domains are fundamentally governed by harmonic superpositions and autoregressive correlations, allowing the pre-trained PFM to act as a universal temporal feature extractor. Moreover, the sequence-wise normalization effectively abstracts away absolute physical time units and amplitude scales, rendering the inputs scale-invariant. Consequently, the pre-trained PFM seamlessly serves as a robust structural prior for the wireless domain, ready to be refined via domain-specific fine-tuning.
@@ III-C1 Input Feature Preparation
We first derive an initial coarse estimate, $\hat{\mathbf{H}}_{\rm P}\in\mathbb{R}^{2N_{\rm t}\times K\times T}$, based on the least squares (LS) method and interpolation, where the real and imaginary parts are concatenated along the antenna dimension. This estimate serves as a noisy approximation of the true CSI. By utilizing the interpolated estimate rather than raw pilot observations, the input tensor to the ViT consistently maintains a full-grid dimension of $2N_{\rm t}\times K\times T$. This unified input representation decouples the network from specific pilot RE indices, thereby ensuring scalability to varying pilot patterns without retraining. To further suppress the noise and leverage its zero-mean property, a despreading operation [37] is adopted to average the estimate across $R_{\rm T}$ and $R_{\rm F}$ neighboring REs in the time and frequency domains, respectively, thereby enhancing feature consistency. The feature maps before and after the despreading operation are concatenated along the time dimension, resulting in a feature tensor of size $2N_{\rm t}\times K\times L_{\rm f}$, where $L_{\rm f}=T+T/R_{\rm T}$. To capture spatial correlations, the features within this tensor are grouped by subcarrier. Specifically, the spatial-time domain feature on each subcarrier is partitioned into $2N_{\rm t}$ patches of length $L_{\rm f}$, which are then linearly projected onto a latent space of dimension $d_{\rm m}$ via a fully-connected (FC) layer. This yields the input embeddings for the subsequent transformer encoder.
@@ III-D Fusion Module
To clarify the dimensional operations in (14), the tensor $\mathbf{U}^{\prime\prime}$ contains $Q$ independent univariate sequences. The fusion residual block treats $Q$ as a parallel batch dimension, independently transforming each sequence’s feature dimension to yield the $Q\times T$ output. Furthermore, “repurposing” the PFM’s output block indicates that the fusion module inherits its identical structural design, e.g., residual connections and activations. However, to accommodate the joint representation, a dimensional adaptation is employed: the input dimension of its first linear layer expands from $d$ to $d+d_{\rm m}$, properly mapping the concatenated hidden states to the $T$-point CSI estimates.
@@ IV-A1 Simulation Settings
The system configurations and simulation settings for channel dataset generation are summarized in Table I. The simulation parameters are randomly sampled from the specified ranges in the table to ensure diverse channel conditions. Each slot consists of $T=14$ OFDM symbols. A total of 8 RBs are allocated, resulting in $K=96$ subcarriers per slot. To emulate realistic propagation environments, the tapped delay line (TDL) channel model compliant with the 3GPP specifications [40] is employed for dataset generation using the MATLAB 5G toolbox. Specifically, the TDL-A, B, and C profiles represent pure non-line-of-sight (NLOS) conditions with Rayleigh fading taps, while the TDL-D profile, which is utilized later for the out-of-distribution zero-shot generalization evaluation, features a line-of-sight (LOS) Rician fading first tap with a fixed K-factor of 13.3 dB. Sparse pilot configurations employing FDM are adopted in the simulations, as depicted in the top-left corner of Fig. 1(a). The “2P” and “4P” patterns denote two different allocations of pilot OFDM symbols under $N_{\rm t}=4$, with $T_{\rm p}=2$ and $T_{\rm p}=4$, respectively, used for different mobility scenarios. These pilot arrangements follow the demodulation RS (DMRS) configuration type 2 defined in 5G NR [41].
In total, 768,000 channel trajectories are generated, with each trajectory comprising 10 consecutive slots. To capture authentic temporal correlations, the channel within each trajectory evolves continuously over time according to the standard 3GPP time-evolution procedure using a Jakes-type Doppler spectrum determined by the designated user speeds and carrier frequency, rather than being independently generated per slot. This results in a dataset of 7,680,000 samples, each comprising a triplet of previous and current channel realizations. The dataset is partitioned into training, validation, and test sets using an 8:1:1 split ratio. It is worth noting that the dataset includes the initial slot configurations where no historical information is available, i.e., $\mathbf{Z}_{\rm his}=\emptyset$. These cold-start samples ensure that the pilot processing network is adequately trained to independently recover the channel when predictive priors are absent.
@@ IV-A3 Baselines
  • LSTM-aided channel estimator (LSTM-aided CE): A lightweight temporal baseline utilizing a long short-term memory (LSTM) model [26, 42] to replace the PFM for providing predictive priors. It comprises approximately 1.26M parameters and is trained using the identical training strategy as the proposed approach. This baseline is specifically included to evaluate the necessity of employing a large-capacity foundation model over lightweight recurrent networks.
@@ IV-B Prediction Performance of PFM
To illustrate the superiority of the PFM and the motivation for its adoption, we compare this model with the state-of-the-art channel prediction method, LLM4CP [26], and the lightweight LSTM baseline [26, 42], as shown in Fig. 5. For a fair comparison, the channel datasets released in [26] are used for training, finetuning, and performance evaluation of the PFM. To explore effective strategies for leveraging the PFM, four training paradigms are examined:
@@
The comparative results, depicted in Fig. 5, lead to several important observations. First, the finetuned PFM substantially outperforms the pre-trained version, indicating that adaptation using wireless channel data effectively enhances model performance in channel acquisition. Second, when comparing models trained from scratch with those finetuned from pre-trained initialization, the results highlight the advantages of leveraging pre-trained weights from the general time-series foundation model [33]. Finally, although freezing the backbone during finetuning achieves competitive performance, updating the backbone parameters yields additional gains and makes the model significantly outperform the LLM4CP baseline, underscoring the importance of jointly refining the entire model to adapt the PFM to wireless channel-specific tasks. These observations substantiate the training strategy proposed in Section III-E for effectively adapting the PFM to wireless channel estimation. Furthermore, the lightweight LSTM baseline significantly lags behind most PFM variants, indicating that conventional lightweight recurrent architectures lack the representational capacity required to capture the complex temporal dependencies and non-linear dynamics inherent in wireless channels.
@@ IV-C Channel Estimation Performance Analysis
Unless otherwise specified (e.g., the slot-wise analysis in Fig. 7), the performance metrics reported in the following evaluations represent the steady-state performance that is averaged over slot indices $i\geq 2$, excluding the cold-start initial slot. This evaluation accurately reflects the steady-state capability and the specific performance gains introduced by the PFM-aided framework.
@@ IV-C1 In-Distribution Performance
We first evaluate the in-distribution performance of the proposed PFM-aided channel estimation scheme (PFM-aided CE) compared to baselines and explore the model’s ability to handle varying configurations, including user speeds and pilot configurations. The channel estimation NMSE performance across different user velocities is presented in Fig. 6. As shown in the figure, PFM-aided CE outperforms all baselines by 1–5 dB in NMSE across all velocities and SNRs under both the “2P” and “4P” pilot configurations. This gain primarily stems from the model’s ability to effectively exploit temporal knowledge from previously estimated CSI. Specifically, at moderate speeds of 30 km/h and 90 km/h, the PFM-aided CE with the sparse “2P” configuration even achieves performance comparable to LMMSE using the denser “4P” configuration, especially in the low-SNR regime. It is worth mentioning that the exceptional results for both the “2P” and “4P” configurations are achieved using the same pre-trained model, validating the architectural scalability of the proposed model and eliminating the overhead of pattern-specific retraining. At high mobility (300 km/h), the temporal correlation of the channel becomes significantly low, leading to severe channel dynamics. As expected, all compared approaches experience performance degradation. However, the proposed estimator still delivers the best performance and achieves even more pronounced gains. These results demonstrate that the PFM can extract complex temporal dependencies even when linear temporal correlation is weak, confirming the ability of the model to handle channel variations with different time granularities.
Moreover, regarding the impact of temporal model capacity, the lightweight LSTM-aided CE significantly underperforms the proposed framework, as shown in Fig. 6. Crucially, under the sparse “2P” configuration, inaccurate LSTM predictions introduce a negative temporal gain, degrading the estimation accuracy even below that of the standalone, non-predictive ViT. This performance penalty highlights that lightweight models fail to provide reliable predictive priors, confirming the essential role of the PFM for high-fidelity channel acquisition.
@@
To evaluate the long-term stability of the recursive inference workflow and address potential concerns regarding exposure bias, i.e., the mismatch between using LMMSE estimates during training and the model’s own outputs during inference, we extend the evaluation trajectory to 15 consecutive slots, well beyond the 10-slot trajectory length used during training. The NMSE performance with regard to the slot index is presented in Fig. 7, where a moderate speed of 90 km/h and SNR of 15 dB are considered. The results show that the proposed estimator maintains stable performance across multiple slots for slot indices $i\geq 2$, clearly outperforming the first slot, which relies solely on the pilot observations for estimation. This observation highlights the effectiveness of the proposed “predict-and-refine” framework. Particularly, PFM-aided CE with “2P” shows virtually the same estimation accuracy as LMMSE with “4P”, despite using only 50% of the pilot overhead, demonstrating its significant potential in reducing pilot overhead.
Furthermore, the absence of error divergence over the 15-slot continuous operation validates the framework’s resilience against error propagation. Unlike pure autoregressive generation where errors can compound, instantaneous pilot observations at each slot serve as anchors to systematically calibrate the predictive priors. Additionally, training the PFM on imperfect LMMSE estimates inherently equips the model to handle noisy historical contexts, ensuring stable adaptation to its own outputs during continuous inference.
Figure 8: NMSE performance of the proposed PFM-aided CE across different spatial correlation levels.
Next, we investigate the effectiveness of the proposed architecture under varying degrees of spatial correlation. It is worth recalling that the PFM branch employs a univariate decomposition strategy, which treats the sequence of each antenna independently. To verify whether the proposed architecture successfully compensates for this lack of spatial awareness during the prediction phase, we evaluate the estimation NMSE across different spatial correlation levels defined in the 3GPP specifications (Low, Medium, Medium-A, and High) [43]. As illustrated in Fig. 8, the estimation accuracy improves as the spatial correlation increases. This trend demonstrates that rather than discarding spatial mutual information across antennas, the proposed “predict-and-refine” framework effectively captures it through the pilot processing ViT. For spatially correlated channels, the fusion module successfully exploits this rich spatial context to calibrate the independent temporal predictive priors provided by the PFM, thereby achieving superior overall channel estimation performance.
@@ IV-C2 Zero-Shot Generalization Study
We further investigate the NMSE performance under an unseen environment using the LOS TDL-D profile, as shown in Fig. 11. This setting introduces a substantial distributional shift in channel correlation and delay structure, as the model was trained exclusively on NLOS scenarios (TDL-A/B/C). Despite this drastic mismatch, the proposed PFM-aided CE maintains competitive performance, remarkably surpassing all the baselines. These results underscore the model’s zero-shot generalization ability, enabling adaptation to structurally different channel environments without finetuning. This capability stems from the time-series forecasting knowledge inherited from the pre-trained PFM. This large-capacity model has been pre-trained across a large-scale time-series corpus, thereby generalizing effectively to out-of-distribution wireless channel sequences, an ability inaccessible to conventional statistical or small-scale AI models.
@@ IV-D End-to-End Bit Error Rate (BER) Performance
Figure 12: End-to-end BER performance comparison of the communication link employing the proposed channel estimator and various baseline methods.
To assess the practical viability of the proposed channel estimator in a communication link, we evaluate its impact on end-to-end BER. We consider a multiple-input single-output OFDM system, and the channel settings follow those presented in Table I. The 5G NR low-density parity-check (LDPC) code with a rate of 466/1024 [44] and a codeword size of 7296 bits is employed to encode the message bits for all data REs. The encoded bits are then modulated using 64-ary quadrature amplitude modulation (64-QAM) and mapped onto the data REs over the resource grid. The data stream is precoded for transmission across $N_{\rm t}=4$ antennas. The transmitter is assumed to perform precoding based on the estimated CSI perfectly fed back from the user. The digital precoder on the $(k,t)$-th RE, $\mathbf{v}_{k,t}\in\mathbb{C}^{4\times 1}$, is obtained based on match filtering [45] using the estimated full CSI $\hat{\mathbf{h}}_{k,t}\in\mathbb{C}^{4\times 1}$ on the $(k,t)$-th RE:
$\mathbf{v}_{k,t}=\frac{\hat{\mathbf{h}}_{k,t}}{\|\hat{\mathbf{h}}_{k,t}\|_{2}}.$ (18)
Hence, the received signal at the $(k,t)$-th RE is given by
@@
where $\mathbf{h}_{k,t}\in\mathbb{C}^{4\times 1}$ represents the true CSI, and $x_{k,t}$ and $n_{k,t}$ denote the transmitted symbol and the AWGN with variance $\sigma^{2}$ at the $(k,t)$-th RE, respectively.
At the receiver side, the received signals are processed using an LMMSE equalizer. To isolate the evaluation to the accuracy of the proposed estimator, the equivalent channel required for equalization is directly reconstructed by the UE using its previously estimated CSI, i.e., $\hat{\mathbf{h}}_{k,t}^{\rm H}\mathbf{v}_{k,t}=\|\hat{\mathbf{h}}_{k,t}\|_{2}$. The equalized signals are then soft-demapped and delivered to the LDPC decoder to recover the transmitted message bits. The BER is evaluated after $10^{4}$ codewords are transmitted.
The BER results obtained using different channel estimators under various mobility conditions are presented in Fig. 12. For reference, the BER performance with perfect CSI at both the transmitter and the receiver is included as an upper bound. Fig. 12(a) presents the results for a user speed of 90 km/h with the “2P” pilot pattern. As shown in the figure, the proposed estimator enables a substantial BER reduction in the high-SNR regime, surpassing all the practical baselines by approximately 4 dB and closely approaching the upper bound. These results highlight the critical role of the accurate channel acquisition achieved by the proposed PFM-aided CE in enhancing the overall system performance with only limited pilot overhead. Fig. 12(b) further evaluates the estimators at a higher user speed of 300 km/h using the “4P” pilot pattern. Although the performance gap between the proposed scheme and the upper bound becomes larger due to increased channel dynamics, the proposed method continues to outperform all baselines by a significant margin, demonstrating strong robustness under high-mobility conditions.